# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ 4       $k\leftarrow k+1$;
$=\mathbb{P}_{\bar{f}}\left(\sum_{i=1}^{n}Z_{i}<-t\right)$
$\leq\exp\left(-\frac{t^{2}/2}{n\mathbb{E}[Z_{1}^{2}]+Kt/3}\right)$
$=\exp\left(-\frac{t^{2}/2}{Kn\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}+Kt/3}\right)$
$=\exp\left(-\frac{3t^{2}}{6Kn\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}+2Kt}\right)$
$=\exp\left(-\frac{3t^{2}}{6KtC^{2}+2Kt}\right)$
$=\exp\left(-\frac{3t}{2K(3C^{2}+1)}\right)$
$=\exp\left(-\frac{3n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\cdot C^{-2}}{16B_{\mathcal{F}}^{2}(3C^{2}+1)}\right)$
$\leq\exp\left(-\frac{3C^{2}\delta^{2}\cdot C^{-2}}{16B_{\mathcal{F}}^{2}(3C^{2}+1)}\right)$
$=\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}(3C^{2}+1)}\right).$
@@
$1-\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
For the other bound, we now set $Z_{i}=|f(X_{i})-\bar{f}(X_{i})|^{2}-\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ for each $i\in[n]$, which are again mean 0 and i.i.d. random variables with $|Z_{1}|\leq K$ for $K=8B_{\mathcal{F}}^{2}$, so that $\sum_{i=1}^{n}Z_{i}=n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}-n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}.$ Then similarly,
$\mathbb{E}[Z_{1}^{2}]$ $\leq\mathbb{E}|f(X_{i})-\bar{f}(X_{i})|^{4}\leq 4B_{\mathcal{F}}^{2}\mathbb{E}|f(X_{i})-\bar{f}(X_{i})|^{2}=K\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}.$
This implies $n\mathbb{E}[Z_{1}^{2}]\leq K\delta^{2}$. Hence using $t=\delta^{2}$,
@@
$\leq\exp\left(-\frac{\delta^{4}/2}{n\mathbb{E}[Z_{1}^{2}]+K\delta^{2}/3}\right)$
$\leq\exp\left(-\frac{\delta^{4}/2}{K\delta^{2}+K\delta^{2}/3}\right)$
$=\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right).$
@@
$1-\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right).$
@@
$1-\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)-\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right),$
we have both $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$.
Lemma 2.6.
Suppose we are testing $H_{0}:\bar{f}=f$ vs $H_{1}:\bar{f}=g$ where $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ for some $C>3$. Then the test $\psi(\vec{Y})=\mathbbm{1}\left(\|\vec{Y}-f(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-g(\vec{X})\|_{2}^{2}\right)$ satisfies
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq 3\exp\left(-L(C,\sigma,B_{\mathcal{F}})\cdot\delta^{2}\right),$
where
$L(C,\sigma,B_{\mathcal{F}})=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{3}{64B_{\mathcal{F}}^{2}},\frac{3}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right\}.$ (2.3)
{toappendix}
@@
We now derive estimates on $\|\eta(\vec{X})\|_{2}$ and $\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ that hold with high probability. Suppose $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. From Lemma 2.5, this condition along with our assumption $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ implies that with probability at least $1-\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)-\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right)$, we have
$n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2},\quad n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}.$ (2.4)
Let $E$ denote the event that both inequalities of (2.4) hold, so
$\mathbb{P}_{\bar{f}}(E^{c})\leq\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
The first inequality in (2.4) implies
@@
The other inequality simply states that $\|\eta(\vec{X})\|_{2}\leq\delta\sqrt{2}$. Thus, the inequalities of (2.4) imply
$\|\eta(\vec{X})\|_{2}\leq\delta\sqrt{2}\leq\frac{\sqrt{2}}{\sqrt{C^{2}-1}}\cdot\|f(\vec{X})-g(\vec{X})\|_{2}$
@@
$-\|f(\vec{X})-g(\vec{X})\|_{2}^{2}+2\|\eta(\vec{X})\|_{2}\|f(\vec{X})-g(\vec{X})\|_{2}$ $\leq\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)\cdot\|f(\vec{X})-g(\vec{X})\|_{2}^{2}$
@@
Using a bound from van1996weak, we note that for any $\rho<0<\tau$,
$\mathbb{P}(N(\rho,\tau^{2})\geq 0)\leq\exp(-\rho^{2}/(2\tau^{2})).$
@@
$\exp\left(-\frac{\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)^{2}n^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{4}}{8n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}}\right)$ $=\exp\left(-\frac{\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)^{2}n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}}{8\sigma^{2}}\right)$
$\leq\exp\left(-\frac{\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)^{2}(C^{2}-1)\delta^{2}}{8\sigma^{2}}\right)$
$=\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)$
Since $\mathbb{P}_{\bar{f}}(\psi=1|\vec{X},E)$ has an upper bound not depending on $X$, it follows (by taking an expectation) that
$\mathbb{P}_{\bar{f}}(\psi=1|E)\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$
@@
$\mathbb{P}_{\bar{f}}(\psi=1)$ $=\mathbb{P}_{\bar{f}}(\psi=1|E)\cdot\mathbb{P}_{\bar{f}}(E)+\mathbb{P}_{\bar{f}}(\psi=1|E^{c})\cdot\mathbb{P}_{\bar{f}}(E^{c})$
@@
$\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
@@
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq\ \exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)$
$\quad\quad\quad+\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
@@
$L(C,\sigma,B_{\mathcal{F}})=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{3}{64B_{\mathcal{F}}^{2}},\frac{3}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right\},$
to conclude that
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)$ $\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)\leq 3\exp\left(-L(C,\sigma,B_{\mathcal{F}})\cdot\delta^{2}\right).$
We now build on the previous lemma by showing the function in a $\delta$-packing closest to our data will also be $\delta$-close (in the $L_{2}(\mathbb{P}_{X})$-distance) to the true function with high probability.
Lemma 2.7.
Let $f_{1},\dots,f_{M}$ be a $\delta$-packing of $\mathcal{F}^{\prime}\subseteq\mathcal{F}$ and suppose $\bar{f}\in\mathcal{F}^{\prime}$. Let $i^{\ast}=\mathop{\mathrm{argmin}}_{i\in[M]}\sum_{j=1}^{n}(Y_{j}-f_{i}(X_{j}))^{2}.$ Then
$\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\delta(C+1)\right)$ $\leq 3M\exp\left(-L(C,\sigma,B_{\mathcal{F}})\cdot n\delta^{2}\right),$
where $L(C,\sigma,B_{\mathcal{F}})$ is defined as in (2.3).
{toappendix}
@@
$\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}+\|f_{i}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}\leq\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}+\delta.$
@@
$\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>\delta(C+1)\right)$ $\leq\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}>\delta\cdot C\right)$
$\leq\mathbb{P}_{\bar{f}}\left(\exists j:\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}>\delta\cdot C\text{ and }\right.$
$\quad\quad\quad\quad\left.\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(\exists j:n\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}^{2}>C^{2}n\delta^{2}\text{ and }\right.$
$\quad\quad\quad\quad\left.\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\right)$
$\leq\sum_{j}\mathbb{P}_{\bar{f}}\left(n\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}^{2}>C^{2}n\delta^{2}\text{ and }\right.$
$\quad\quad\quad\quad\left.\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\right)$
$\leq 3M\exp\left(-L(C,\sigma,B_{\mathcal{F}})\cdot n\delta^{2}\right).$ (2.5)
where we used Lemma 2.6 in (2.5) with $\sqrt{n}\delta$ in place of $\delta$.
Remark 2.8.
Lemma 2.6 and Lemma 2.7 can be generalized to our sub-Gaussian setting where each $\xi_{i}\sim\mathrm{SG}(0,\sigma^{2})$. We supply the details in the appendix.
{toappendix}
@@
Moreover, the variance proxy can be shown to be $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ using the definition of a sub-Gaussian random variable. The proof of Lemma 2.6 then defines the event $E$ as when the inequalities in (2.4) hold, and $E$ satisfies
$\mathbb{P}_{\bar{f}}(E^{c})\leq\exp\left(-\frac{3\delta^{2}}{64B_{\mathcal{F}}^{2}}\right)+\exp\left(-\frac{3\delta^{2}}{16B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
@@
$\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$
@@
and variance proxy $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$. A sub-Gaussian random variable $Z$ with mean $\rho$ and variance proxy $\tau^{2}$ can be shown to satisfy $P(Z>t+\rho)\leq\exp(-t^{2}/2\tau^{2})$ for any $t>0$ (boucheron2013, Section 2.3), and if $\rho<0$, we can take $t=-\rho$ and obtain a bound of $P(Z>0)\leq\exp(-\rho^{2}/2\tau^{2})$. Applying this result, as in the previous lemma, we conclude that conditional on $\vec{X}$ and the event $E$, $\psi=1$ holds with probability bounded by
$\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$

### v5

@@ Appendix B Proofs for Section 3
$\leq\exp\left(-\frac{t^{2}}{16nL\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}+4Lt}\right)$
$=\exp\left(-\frac{t^{2}}{16C^{2}Lt+4Lt}\right)$
$=\exp\left(-\frac{t}{4L(C^{2}+1)}\right)$
$=\exp\left(-\frac{C^{-2}n\|f-g\|_{L_{2}(\mathbb{P}_{X})}}{4L(C^{2}+1)}\right)$
$\leq\exp\left(-\frac{\delta^{2}}{4L(C^{2}+1)}\right).$
@@
$1-\exp\left(-\frac{\delta^{2}}{4L(C^{2}+1)}\right).$
For the other bound, we now set $Z_{i}=|f(X_{i})-\bar{f}(X_{i})|$ for each $i\in[n]$, so that we deduce the same bound (B.4) with $\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}$ and $\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}$ instead. Setting $t=\delta^{2}$ and repeatedly using $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$, we have
@@
$\leq\exp\left(-\frac{t^{2}}{16nL\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}+4Lt}\right)$
$\leq\exp\left(-\frac{t^{2}}{16L\delta^{2}+4Lt}\right)$
$=\exp\left(-\frac{\delta^{2}}{20L}\right).$
@@
$1-\exp\left(-\frac{\delta^{2}}{20L}\right).$
@@
$1-\exp\left(-\frac{\delta^{2}}{20L}\right)-\exp\left(-\frac{\delta^{2}}{4L(C^{2}+1)}\right),$
we have both $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$. ∎
@@
We now derive estimates on $\|\eta(\vec{X})\|_{2}$ and $\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ that hold with high probability. Suppose $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. From Lemma 3.7, this condition along with our assumption $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ implies that with probability at least $1-\exp\left(-\frac{\delta^{2}}{20L}\right)-\exp\left(-\frac{\delta^{2}}{4L(C^{2}+1)}\right)$, we have
$n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2},\quad n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}.$ (B.5)
Let $E$ denote the event that both inequalities of (B.5) hold, so
$\mathbb{P}_{\bar{f}}(E^{c})\leq\exp\left(-\frac{\delta^{2}}{20L}\right)+\exp\left(-\frac{\delta^{2}}{4L(C^{2}+1)}\right).$
The first inequality in (B.5) implies
@@
The other inequality simply states that $\|\eta(\vec{X})\|_{2}\leq\delta\sqrt{2}$. Thus, the inequalities of (B.5) imply
$\|\eta(\vec{X})\|_{2}\leq\delta\sqrt{2}\leq\frac{\sqrt{2}}{\sqrt{C^{2}-1}}\cdot\|f(\vec{X})-g(\vec{X})\|_{2}.$
@@
$\leq\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)\cdot\|f(\vec{X})-g(\vec{X})\|_{2}^{2}$
@@
Using a bound from Van Der Vaart and Wellner [1996, Section 2.2.1], we note that for any $\rho<0<\tau$,
$\mathbb{P}(\mathcal{N}(\rho,\tau^{2})\geq 0)\leq\exp(-\rho^{2}/(2\tau^{2})).$
@@
$\exp\bigg{(}-\frac{\big{(}-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\big{)}^{2}n^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{4}}{8n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}}\bigg{)}$
$=\exp\bigg{(}-\frac{\big{(}-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\big{)}^{2}n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}}{8\sigma^{2}}\bigg{)}$
$\leq\exp\bigg{(}-\frac{\left(-1+\frac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)^{2}(C^{2}-1)\delta^{2}}{8\sigma^{2}}\bigg{)}$
$=\exp\bigg{(}-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\bigg{)}.$
Since $\mathbb{P}_{\bar{f}}(\psi=1|\vec{X},E)$ has an upper bound not depending on $\vec{X}$, it follows (by taking an expectation) that
$\mathbb{P}_{\bar{f}}(\psi=1|E)\leq\exp\bigg{(}-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\bigg{)}.$
@@
$\mathbb{P}_{\bar{f}}(\psi=1)$
$=\mathbb{P}_{\bar{f}}(\psi=1|E)\cdot\mathbb{P}_{\bar{f}}(E)+\mathbb{P}_{\bar{f}}(\psi=1|E^{c})\cdot\mathbb{P}_{\bar{f}}(E^{c})$
@@
$\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)+\exp\left(-\frac{\delta^{2}}{20L}\right)$
$\quad\quad+\exp\left(-\frac{\delta^{2}}{4L(C^{2}+1)}\right).$
@@
$\sup_{\bar{f}:n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=1)\vee\sup_{\bar{f}:n\|g-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}}\mathbb{P}_{\bar{f}}(\psi=0)$
$\leq\exp\left(-\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right)+\exp\left(-\frac{\delta^{2}}{20L}\right)$
$\quad\quad+\exp\left(-\frac{\delta^{2}}{4L(C^{2}+1)}\right).$
@@
$\Lambda(C,\sigma,L)=\min\left\{\frac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}}{8\sigma^{2}},\frac{1}{20L},\frac{1}{4L(C^{2}+1)}\right\},$
and replace the preceding upper bound by $3\exp\left(-\Lambda(C,\sigma,L)\cdot\delta^{2}\right).$
@@
$\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}+\|f_{i}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}$
$\leq\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}+\delta.$
@@
$\mathbb{P}_{\bar{f}}\big{(}\|f_{i^{\ast}}-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}>(C+1)\delta\big{)}\leq\mathbb{P}_{\bar{f}}\left(\|f_{i^{\ast}}-f_{i}\|_{L_{2}(\mathbb{P}_{X})}>C\delta\right)$
$\leq\mathbb{P}_{\bar{f}}\Big{(}\exists j:\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}>C\delta\text{ and }$
$\qquad\qquad\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\Big{)}$
$=\mathbb{P}_{\bar{f}}\Big{(}\exists j:n\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}^{2}>C^{2}n\delta^{2}\text{ and }$
$\qquad\qquad\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\Big{)}$
$\leq\sum_{j}\mathbb{P}_{\bar{f}}\Big{(}n\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}^{2}>C^{2}n\delta^{2}\text{ and }$
$\qquad\qquad\|\vec{Y}-f_{i}(\vec{X})\|_{2}^{2}\geq\|\vec{Y}-f_{j}(\vec{X})\|_{2}^{2}\Big{)}$
$\leq 3M\exp\big{(}-\Lambda(C,\sigma,L)\cdot n\delta^{2}\big{)}.$ (B.6)
where we used Lemma 3.8 in (B.6) with $\sqrt{n}\delta$ in place of $\delta$. ∎
@@
Moreover, the variance proxy can be shown to be $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ using the definition of a sub-Gaussian random variable. The proof of Lemma 3.8 then defines the event $E$ as when the inequalities in (B.5) hold, and $E$ satisfies
$\mathbb{P}_{\bar{f}}(E^{c})\leq\exp\left(-\tfrac{\delta^{2}}{20L}\right)+\exp\left(-\tfrac{\delta^{2}}{4L(C^{2}+1)}\right).$
@@
$\left(-1+\tfrac{2\sqrt{2}}{\sqrt{C^{2}-1}}\right)\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$
@@
and variance proxy $4n\sigma^{2}\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$. A sub-Gaussian random variable $Z$ with mean $\rho$ and variance proxy $\tau^{2}$ can be shown to satisfy $P(Z>t+\rho)\leq\exp(-t^{2}/2\tau^{2})$ for any $t>0$ [Boucheron et al., 2013, Section 2.3], and if $\rho<0$, we can take $t=-\rho$ and obtain a bound of $P(Z>0)\leq\exp(-\rho^{2}/2\tau^{2})$. Applying this result, as in the previous lemma, we conclude that conditional on $\vec{X}$ and the event $E$, $\psi=1$ holds with probability bounded by
$\exp\left(-\tfrac{\left(\sqrt{C^{2}-1}-2\sqrt{2}\right)^{2}\delta^{2}}{8\sigma^{2}}\right).$